OpenStudy (anonymous):

sqrt 2 + sqrt 3 + sqrt 2

OpenStudy (kkutie7):

sqrt(2+3+2)= sqrt(7)

OpenStudy (anonymous):

im suppose to simplify

OpenStudy (kkutie7):

yeah this is simplifying

geerky42 (geerky42):

\(\sqrt{2}+\sqrt{3}+\sqrt{2}\neq\sqrt{7}\) Because \(\sqrt{a}+\sqrt{b}\neq\sqrt{a+b}\)

geerky42 (geerky42):

You can simplify that part: \(\sqrt{2}+\sqrt{2}\)